Suppose 8 quarts of a solution that is 45% antifreeze is mixed with 12 quarts of a solution that is
33% antifreeze. How many quarts of antifreeze are in the resulting mixture? What percent of the resulting mixture is antifreeze?

Answer by richwmiller(17219)

.45*8+.33*12=20x
3.60+3.96=20x
7.56=20x
x=7.56/20
x=.378
x=37.8%
a+b=20,
0.33*a+0.45*b=0.378*20
a=20-b
0.33*(20-b)+0.45*b=7.56
6.6-0.33b+0.45*b=7.56
0.12*b=0.96
b=8.0
a=20-b
a=12.0 qt. at 33%
b=8.0 qt. at 45%
check
0.33*12.0+0.45*8.0=0.378*20
3.96+3.6=7.56
7.56=7.56
ok